OpenStudy (anonymous):

Hello, this is my first time using OpenStudy aside from browsing through how to do certain problems. Here's one I'm having trouble on: A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(3t). Find the acceleration of the particle when the velocity is first zero. I tried getting the second derivative and plugging in zero, but that doesn't seem to be right. Could somebody help?

OpenStudy (solomonzelman):

The acceleration is the second derivative

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle s(t)=t\ln(3t) }\) this is your position function, correct?

OpenStudy (anonymous):

That is correct.

OpenStudy (solomonzelman):

In general, the derivative of the function is the function's slope, and this derivative will tell us how fast (or how "steep") the function is going at every point. And so, the derivative (the slope) of the position tells us how fast this position will move. So, wouldn't you agree that the derivative of the position, is the velocity?

OpenStudy (anonymous):

Yes, and the derivative of velocity is the acceleration.

OpenStudy (solomonzelman):

can you differentiate the s'(t)? (to find the general expression for velocity of the particle)?

OpenStudy (solomonzelman):

I mean find the s'(t) (not differentiate s'(t))....

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle s(t)=t\ln(3t) }\)

OpenStudy (solomonzelman):

use the product rule to differentiate this

OpenStudy (anonymous):

ln(3t) + 1, correct?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle s(t)=t\ln(3t) }\) \(\color{#000000 }{ \displaystyle s(t)=t\ln(t)+t\ln(3) }\) \(\color{#000000 }{ \displaystyle s'(t)=1\cdot\ln(t)+t\cdot \frac{1}{t}+\ln(3) }\) \(\color{#000000 }{ \displaystyle s'(t)=\ln(t)+1+\ln(3) }\) yes, your derivative is corret

OpenStudy (solomonzelman):

Now, you are asked to find the acceleration at the points where the particle's velocity is zero.

OpenStudy (anonymous):

Do I plug in zero, then take the derivative of that?

OpenStudy (anonymous):

Wait, set it to zero rather.

OpenStudy (anonymous):

shouldn't the domain be t>0? as t ->0+, ln(t) -> - infinity

OpenStudy (anonymous):

I don't think so. That's the domain I was given.

OpenStudy (anonymous):

so what's the position at t = 0? doesn't make sense.

OpenStudy (anonymous):

sorry, I don't mean to distract...

OpenStudy (anonymous):

@solomonzelman Any insight you'd be willing to give? I'm rather confused.

OpenStudy (anonymous):

Of course aside from guiding me right up to the point where I was initially confused--my initial steps were shown to be correct.

OpenStudy (solomonzelman):

I had technological difficulties... apologize

OpenStudy (solomonzelman):

pgpilot326, it is asking for the points where velocity is 0, not s'(0)

OpenStudy (anonymous):

Don't sweat it! I really appreciate your guidance.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle s(t)=t\ln(3t) }\) Find the acceleration of the particle when the velocity is first zero. \(\color{#000000 }{ \displaystyle s'(t)=1\cdot \ln(3t) +t\times \frac{1}{3t}\times 3 = \ln(3t)+1 }\) You need to know the point(s) where velocity is =0. \(\color{#000000 }{ \displaystyle 0=1\cdot \ln(3t) +t\times \frac{1}{3t}\times 3 = \ln(3t)+1 }\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 0= \ln(3t)+1 }\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle -1= \ln(3t) }\) \(\color{#000000 }{ \displaystyle -1= \log_e(3t) }\) \(\color{#000000 }{ \displaystyle e^{-1}= 3t }\) go on...

OpenStudy (solomonzelman):

the solution to that (to s'(t)=0), is going to give you the point where the velocity of the particle is zero (say that you find that \(x=\beta\). (Just making it up) Keep that in mind, because you want to find the acceleration at the point \(x=\beta\) \( (\)since the question asks for acceleration at point(s) where velocity is zero\( ) \), so you need to find the second derivative \(s''(t)\) which gives you the acceleration for \(s(t)\) at every point, and then to plug in \(s''(\beta)\) to find the acceleration at a point \(x=\beta\) \( (\) Which is the acceleration wanted - the acceleration of the particle at the point where the velocity is equal to zero\( )\). This is verbal, and if you want I can give you an example similar to this.

OpenStudy (anonymous):

Alright so I ended up with t = 1/3e ; however, that's not a choice. Forgive me. I completely understand the conceptual aspects you've pointed out, but this isn't a choice. Hm.

OpenStudy (solomonzelman):

1/(3e) is just a point where velocity is 0.

OpenStudy (solomonzelman):

yiou will need acceleration at this point, so find the second derivative of the function s(t), and evaluate it at t=1/(3e)

OpenStudy (solomonzelman):

So, you need f''(1/(3e))

OpenStudy (anonymous):

Yes you're right!! Sorry haha I got it now!

OpenStudy (anonymous):

Thanks so much. That was dumb of me.

OpenStudy (solomonzelman):

No problem.... unlike many users you preform the work correctly, and I enjoy that! Ty and yw ;)

OpenStudy (irishboy123):

welcome @Jerobrien !!!

OpenStudy (solomonzelman):

Similar example (just in case): A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = 2t•In(6t). Find; \(\small\color{blue}{(a)}\) Acceleration of the particle when the velocity is =0. \(\small\color{blue}{(b)}\) Acceleration of the particle when the velocity is =2. \(\small\color{blue}{(c)}\) Prove that the particle's velocity always increases. \(\color{#000000 }{ \displaystyle s(t)=2t\ln(6t) }\) Velocity/Derivative \(\color{#000000 }{ \displaystyle s'(t)=2\ln(6t)+2 }\) Velocity is 0, \(\color{#000000 }{ \displaystyle 0=2\ln(6t)+2 }\) \(\color{#000000 }{ \displaystyle -2=\ln(6t) }\) \(\color{#000000 }{ \displaystyle e^{-2}=6t }\) \(\color{#000000 }{ \displaystyle t=1/(6e) }\) Velocity is 2, \(\color{#000000 }{ \displaystyle 2=2\ln(6t)+2 }\) \(\color{#000000 }{ \displaystyle 0=\ln(6t) }\) \(\color{#000000 }{ \displaystyle t=1/6}\) Acceleration/2nd Derivative \(\color{#000000 }{ \displaystyle s''(t)=2/t }\) Acceleration at the point where velocity is 0, \(\color{#000000 }{ \displaystyle s''(1/(6e))=2/(1/(6e)) }\) \(\color{#000000 }{ \displaystyle s''(1/(6e))=12e }\) ANSWER \(\small\color{blue}{(a)}\) Acceleration at the point where velocity is 2, \(\color{#000000 }{ \displaystyle s''(1/6)=2/(1/6) }\) \(\color{#000000 }{ \displaystyle s''(1/(6e))=12 }\) ANSWER \(\small\color{blue}{(b)}\) The time is always positive, and for all positive t, \(\color{#000000 }{ \displaystyle s''(t)=2/t }\) is going to be positive. And positive acceleration, by definition is the increase in velocity. ((accel. is always pos. = veloc. always increases)) ANSWER \(\small\color{blue}{(c)}\)

OpenStudy (solomonzelman):

Although velocity is increasing, the acceleration is slowing down; \(\color{#000000 }{ \displaystyle \frac{ds }{dt} \left[\frac{2}{t}\right] =-\frac{2}{t^2}}\) And the derivative of accelration is always negative, therefore the acceleration is decreasing.

OpenStudy (anonymous):

@Solomonzelman thanks so much again! I went through that example of yours and I do understand it completely now. Also, thanks @irishboy123 :)

OpenStudy (solomonzelman):

Anytime! Enjoy!